OpenStudy (anonymous):

Find the slope of the following line. Leave answers in fractional form. 3x - 2y = 4

OpenStudy (anonymous):

2y=3x-4 y=(3/2)x-4 on comparing with the equation y=mx+c m=3/2

OpenStudy (anonymous):

y=(3/2)x-2